A random variable X has Poisson distribution with mean 2. Then P(X > 1.5) equals ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
P(x = k) =e−λ(λk/k!)
P(x ≥ 2) = 1 – P(x = 0) – P(x = 1)
= 1 – e-λ – e-λ(λ/1!)
=1-(3/e2)
Thanks and Regards
Shaik Aasif
askIITians faculty